# Aerial Active STAR-RIS-assisted Satellite-Terrestrial Covert Communications

### v1

@@ III-B Problem Formulation
As outlined earlier, we define the fairness index of the channel capacity according to Jain’s index to guarantee the fairness among all ground users, which can be represented as follows [23]:
@@
Note that the specific challenges associated with the formulated ASCCOP are listed as follows.
@@
Conventional methods typically struggle with this because they require knowledge of future states, which are often uncertain and not directly observable in real-time. In this case, DRL has shown promise in addressing dynamic and uncertain environments in wireless communications [24], [25]. Therefore, we transform the ASCCOP in the context of DRL and design an effective DRL-based algorithm to solve it.
@@ IV-A MDP Formulation
The dynamic decision process in the proposed ASCCOP can be naturally modeled as an MDP, which can be described as a $5$-tuple $\Omega=<\mathcal{S},\mathcal{A},R,P,\gamma>$ [26]. At each time slot $n$, the agent receives the state $\mathbf{s}[n]$ from the environment and selects an action $\mathbf{a}[n]\in\mathcal{A}$ based on its policy $\pi(\mathbf{a}|\mathbf{s})$. Then, the next state is generated based on the transition probability $P(\mathbf{s^{\prime}}|\mathbf{s},\mathbf{a})$, and the agent receives the reward $r[n]=R(\mathbf{s}[n],\mathbf{a}[n])$ from the environment as a feedback. Moreover, $\gamma\in(0,1]$ is the discount factor that balances immediate and future rewards. Accordingly, the details of each element in the considered ASCCOP are explained as follows.
@@ IV-A1 Action
$\mathbf{a}[n]=\left\{\mathbf{v}_{r}[n],\beta_{1}[n],\dots,\beta_{M}[n],\phi_{1}[n],\dots,\phi_{M}[n]\right\}.$ (19)
@@ IV-A2 State
$r_{a1}[n-1],\dots,r_{aK}[n-1],r[n-1]\}.$
@@ IV-A3 Reward
$r[n]=J[n]\sum_{k=1}^{K}r_{ak}[n]+r_{pc}+r_{pr}+r_{pp},$ (21)
where $r_{pc}$ is the penalty when the minimal DEP falls below the required threshold. Moreover, $r_{pr}$ is the penalty when the AASTAR-RIS exceeds the maximum power consumption and $r_{pp}$ is the penalty when its trajectory adjustments violate the movement or area constraints.
@@ IV-B Conventional Deterministic Policy Gradient Algorithm
The DPG algorithm is a classic reinforcement learning method that combines both policy-based and value-based methods [27], which enables simultaneous action generation and evaluation, thus improving training efficiency and decision-making in complex environments. Specifically, the main idea of the DPG algorithm primarily relies on policy evaluation and policy improvement to optimize decision-making processes. Let $\{\mathbf{s}[n],\mathbf{a}[n],\mathbf{s}[n+1],r[n]\}_{n\geq 0}\sim\pi_{\theta}$ be a trajectory sampled from the parameterized policy $\pi_{\theta}$, then the DPG algorithm aims to find an optimal policy $\pi_{\theta}^{*}$ that maximizes the expected cumulative return, which can be expressed as follows:
@@ IV-B2 Policy Improvement
The policy gradient theorem provides a foundation for policy improvement and this process can be expressed as follows [27]:
@@ IV-C Generative Deterministic Policy Gradient Algorithm
IV-C1 GDM for Policy Representation
When addressing the proposed ASCCOP, the conventional DPG algorithm faces two primary challenges. First, this method inherently struggles to explore the complex, high-dimensional and multimodal state-action spaces due to its reliance on unimodal Gaussian distribution [28]. Second, the frequent penalties imposed by covert constraints further exacerbate this issue by discouraging the agent from effectively leveraging the rare positive reinforcement signals.
Compared to fully-connected deep neural networks adopted in the conventional DPG algorithm, GDM provides certain distinct advantages. Specifically, GDM excels in generating the diverse and high-quality samples, which is particularly beneficial for exploring the vast and complex state-action spaces [29]. Moreover, the superior feature representation capability of GDM enables a deeper understanding of the environmental state, which is crucial for effective decision-making in complex scenarios.
According to [29], the basic process of GDM includes the forward process and reverse process. Specifically, the forward process of GDM incrementally corrupts a data sample by adding Gaussian noise over a series of time steps, thereby transforming its distribution into one that approximates pure Gaussian noise. More formally, this forward procedure constitutes a Markov chain, which can be defined as follows:
where $T$ is the total number of diffusion steps, and $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathscr{N}(\mathbf{x}_{t};\sqrt{1-\nu_{t}}\mathbf{x}_{t-1},\nu_{t}\mathbf{I})$ represents the transition kernel of the forward process. Moreover, $\nu_{t}\in(0,1)$ is a diffusion weight to control the speed of diffusion. After applying a series of derivations, the joint transition over all time steps can be marginalized to yield a closed-form expression for the distribution of $\mathbf{x}_{t}$ conditioned on $\mathbf{x}_{0}$, which can be computed as follows [30]:
where $\mu_{\theta}(\mathbf{x}_{t},t)$ and $\Sigma_{\theta}(\mathbf{x}_{t},t)$ represent the mean and the covariance matrix of the predicting noise at time step $t$, respectively. To incorporate specific semantic attributes into the reverse process, the conditional GDM can provide additional information to guide each denoising step [31]. Specifically, the reverse transition chain can be further expressed as follows:
IV-C2 Action Gradient for Policy Improvement
As shown in Eq. (30), GDM merely learns to approximate the distribution of training samples produced by the current policy $\pi_{\theta}$. As such, it does not intrinsically offer a mechanism to optimize the policy beyond the scope of the data on which it is trained. Different from the conventional DPG algorithm, direct policy improvement through the policy gradient theorem is not feasible in this setting, as the diffusion-based policy is not parameterized straightforwardly but is instead learned via a stochastic process. Therefore, we adopt an action gradient mechanism to enable policy improvement, which is shown in Fig. 2. Specifically, a given state-action pair $(\mathbf{s},\mathbf{a})$ can be converted into a better state-action pair by the gradient ascent method, which can be represented as follows [30]:
@@ 4       $\mathbf{z}\sim\mathscr{N}(\mathbf{0},\mathbf{I})$ if $t>1$ else $\mathbf{z}=\mathbf{0}$;
1 Initialize GDM network parameters $\theta$, critic network parameters $\phi_{1}$ and $\phi_{2}$ as well as their corresponding target network parameters ${t\theta}$, $t\phi_{1}$ and $t\phi_{2}$;
@@ 11                   Sample a batch from $\mathcal{D}$;
Figure 3: The overall architecture of the proposed GDPG algorithm to solve the ASCCOP.
@@ IV-D Overall Structure and Complexity Analysis
Similar to approaches used in [32], [33], we employ a target GDM network $\pi_{t\theta}$, and a double critic networks $Q_{\phi_{1}}$ and $Q_{\phi_{2}}$ alongside their corresponding target critic networks $Q_{t\phi_{1}}$ and $Q_{t\phi_{2}}$ to avoid abrupt distribution shifts and address Q-value overestimation. Specifically, the loss function of the critic network can be expressed as follows:
@@ IV-D1 Training Phase
Let $G$, $B$, $|\bm{\phi}|$ and $|\bm{\theta}|$ represent the number of training episodes, batch size, the number of critic network parameters, and the number of GDM network parameters, respectively. Then, the total computational complexity during the training phase is $\mathcal{O}(GN(B+(4+T)|\bm{\theta}|+6|\bm{\phi}|))$. Specifically, the computational complexity of the training phase is summarized as follows [30]:
  • Network Initialization: This process involves initializing two critic networks, a GDM network, and their target networks, thereby increasing the overhead $\mathcal{O}(4|\bm{\phi}|+2|\bm{\theta}|)$.
  • Action Selection: GDM-based action generation at each step needs to performs $T$ denoising steps, thereby introducing the overhead $\mathcal{O}(GNT|\bm{\theta}|)$ during the entire training phase.
  • Network Updates: The overhead of the network update process includes several components. The parameter update for the two critic networks incurs the overhead $\mathcal{O}(2GN|\bm{\phi}|)$. Moreover, action gradient improvements add the overhead $\mathcal{O}(GNB)$. Additionally, the GDM network update introduces the overhead $\mathcal{O}(GNT|\bm{\theta}|)$. Finally, the soft update operation results in the overhead $\mathcal{O}(GN|\bm{\phi}|+2GN|\bm{\theta}|)$. As a result, the overall complexity is $\mathcal{O}(GN(B+2|\bm{\theta}|+4|\bm{\phi}|))$.
@@ IV-D2 Execution Phase
Figure 4: Architecture of GDM network with Mish activation function [34] for non-monotonic gradient flow, and sinusoidal position embedding for efficient encoding of temporal information in the reverse process.
@@ V-A Simulation Setup
V-A1 Environment and Model Details
In this work, we consider the service area as a square region with 200 m by 200 m. In this region, 5 ground users are served by an LAP, operating at $H=100$ m, and equipped with a $4\times 4$ active STAR-RIS array to facilitate the wireless covert transmission between a GEO satellite and the ground users. For the proposed GDPG algorithm, we utilize the Adam optimizer with a learning rate $3\times 10^{-4}$ for both the GDM network and the critic networks, while setting $\eta_{a}=3\times 10^{-2}$ for the action gradient. The specific GDM network structure is exhibited in Fig. 4 and other key parameters for the environment and algorithm are summarized in Table II.
@@ V-A1 Environment and Model Details
$\sigma_{r}^{2}$, $\sigma_{k}^{2}$, $\sigma_{w}^{2}$174 dBm/Hz $\rho$ 3 dB
Algorithm $\gamma$ 0.95 $T$ 10
@@ V-A2 Benchmarks
  • State-of-the-Art DRL Algorithms: The deep deterministic policy gradient (DDPG) and twin delayed DDPG (TD3) are the earliest actor-critic algorithms designed for continuous action spaces [32]. Moreover, soft actor-critic (SAC) is based on the maximum entropy reinforcement learning framework and also serves as a baseline [33].
In order to make a fair performance comparison with benchmarks, the neural network architectures of all DRL-based benchmarks comprise two hidden layers for both the actor and critic networks, with each layer containing 256 neurons. In addition, the learning rate is consistent with the proposed GDPG algorithm.
@@ V-B2 Comparison with State-of-the-Art DRL Algorithms
Fig. 6(a) illustrates the reward per episode of the proposed GDPG algorithm in comparison to other state-of-the-art DRL algorithms during the training process. As can be seen, the proposed GDPG algorithm consistently outperforms the other algorithms with a steady increase in reward over the training time. The success of the GDPG algorithm lies in the use of the GDM, which improves the exploration and policy optimization by generating the diverse and high-quality samples through its denoising process, thereby resulting in higher overall rewards, especially in the complex satellite-terrestrial covert communication scenarios. Moreover, DDPG and TD3 demonstrate initial improvements in rewards during the early stages of training, but their progress slows and eventually plateaus. This indicates that although they achieve early gains, their performance is constrained by limited exploration, thereby leading to suboptimal policy. Finally, SAC performs better than DDPG and TD3 with a more consistent increase in rewards. However, it still has lower performance than that of the GDPG algorithm, thereby illustrating that SAC lacks the same level of exploration efficiency and optimization capabilities provided by the proposed GDPG algorithm.
Fig. 6(b) shows the mean sum channel capacity per step of the proposed GDPG algorithm in comparison to other state-of-the-art DRL algorithms. The performance achieved by the proposed GDPG algorithm is notably higher than that of other algorithms. This indicates that the GDM-enabled DRL algorithm is more effective in leveraging its representation and exploration capabilities in the large state and action spaces. Another important metric is the fairness index, which is presented in Fig. 6(c). As can be seen, the GDPG algorithm maintains the highest fairness index across all algorithms. While SAC converges faster and experiences greater fluctuations in terms of the fairness index when compared to other algorithms, the GDPG algorithm achieves the fastest convergence and consistently maintains a higher fairness index, owing to its denoising process and more effective exploration of the action space.

### v2

@@ III-B Problem Formulation
As outlined earlier, we define the fairness index of the channel capacity according to Jain’s index to guarantee the fairness among all ground users, which can be represented as follows [Sediq2013]:
@@
III-C Problem Analysis
The formulated ASCCOP presents some specific challenges, which are analyzed as follows.
  • Complex Trade-offs among Objectives and Constraints: Sum channel capacity, fairness among ground users and covertness constraint have inherent trade-offs. Specifically, maximizing the sum channel capacity tends to prioritize users with favorable channel conditions, thereby compromising fairness among ground users. Meanwhile, enhancing the sum channel capacity often necessitates increasing the transmit power of the AASATR-RIS, which elevates the probability of detection by the Warden. Furthermore, enforcing fairness for ground users in disadvantaged locations typically requires increasing transmission power or maneuvering the AASTAR-RIS closer to those users. However, such behaviors cause a simultaneous decrease in covertness performance when the Warden is near such users.
@@ III-C Problem Analysis
Conventional methods, e.g., convex optimization and dynamic programming, typically struggle with this because they require knowledge of future states, which are often uncertain and not directly observable in real-time. In this case, DRL has shown promise in addressing dynamic and uncertain environments in wireless communications [Li2024a], [Li2024]. Therefore, we transform the ASCCOP in the context of DRL and design an effective DRL-based algorithm to solve it.
@@ IV-A MDP Formulation
The dynamic decision process in the proposed ASCCOP can be naturally modeled as an MDP, which can be described as a $5$-tuple $\Omega=<\mathcal{S},\mathcal{A},R,P,\gamma>$ [Yimengjie2025]. At each time slot $n$, the agent receives the state $\mathbf{s}[n]$ from the environment and selects an action $\mathbf{a}[n]\in\mathcal{A}$ based on its policy $\pi(\mathbf{a}|\mathbf{s})$. Then, the next state is generated based on the transition probability $P(\mathbf{s^{\prime}}|\mathbf{s},\mathbf{a})$, and the agent receives the reward $r[n]=R(\mathbf{s}[n],\mathbf{a}[n])$ from the environment as a feedback. Moreover, $\gamma\in(0,1]$ is the discount factor that balances immediate and future rewards. Accordingly, the details of each element in the considered ASCCOP are explained as follows.
@@ IV-A1 Action
$\mathbf{a}[n]=\left\{\mathbf{v}_{r}[n],\beta_{1}[n],\dots,\beta_{M}[n],\phi_{1}[n],\dots,\phi_{M}[n]\right\},$ (19)
where the dimension of the action space is $2+2M$.
@@ IV-A2 State
$r_{a1}[n-1],\dots,r_{aK}[n-1],r[n-1]\},$
where the dimension of the state space is $2M(K+1)+3K+9$.
@@ IV-A3 Reward
$r[n]$ $=J[n]\sum_{k=1}^{K}r_{ak}[n]$ (21)
$\qquad-r_{pc}\cdot\mathbb{I}\left\{\iota[n]\geq\min\Bigg(\rho\hat{\sigma}_{w}^{2}-\frac{\hat{\sigma}_{w}^{2}}{\rho},\frac{\left(\rho^{2\varepsilon}-1\right)\hat{\sigma}_{w}^{2}}{\rho}\Bigg)\right\}$
$\qquad-r_{pr}\cdot\mathbb{I}\left\{\left\|\bm{\Phi}[n]\mathbf{h}_{ar}[n]\right\|^{2}+\sigma_{r}^{2}\left\|\bm{\Phi}[n]\right\|^{2}\geq P_{\max}^{\mathrm{active}}\right\}$
$\qquad-r_{pp}\cdot\mathbb{I}\left\{\mathbf{q}_{r}[n+1]\notin[X_{\min},X_{\max}]\times[Y_{\min},Y_{\max}]\right\},$
where $r_{pc}$ imposes a fixed penalty when the covert constraint is violated as determined by the theoretical detection error bound, $r_{pr}$ penalizes any violation of the AASTAR-RIS instantaneous power budget, and $r_{pp}$ penalizes trajectory violations beyond the operating boundaries. Moreover, the indicator function $\mathbb{I}\{\cdot\}$ returns 1 when the specified condition is satisfied and 0 otherwise. Note that the continuous action space with non-convex feasibility regions makes explicit illegal action analytically intractable, which motivates us to adopt the penalty-based reward design method rather than action masking.
@@ IV-B Conventional Deterministic Policy Gradient Algorithm
The DPG algorithm is a classic reinforcement learning method that combines both policy-based and value-based methods [Silver2014], which enables simultaneous action generation and evaluation, thus improving training efficiency and decision-making in complex environments. Specifically, the main idea of the DPG algorithm primarily relies on policy evaluation and policy improvement to optimize decision-making processes. Let $\{\mathbf{s}[n],\mathbf{a}[n],\mathbf{s}[n+1],r[n]\}_{n\geq 0}\sim\pi_{\theta}$ be a trajectory sampled from the parameterized policy $\pi_{\theta}$, then the DPG algorithm aims to find an optimal policy $\pi_{\theta}^{*}$ that maximizes the expected cumulative return, which can be expressed as follows:
@@ IV-B2 Policy Improvement
The policy gradient theorem provides a foundation for policy improvement and this process can be expressed as follows [Silver2014]:
@@ IV-C Generative Deterministic Policy Gradient Algorithm
IV-C1 Motivation of Adopting GDM
When addressing the proposed ASCCOP, the conventional DPG algorithm faces two primary challenges. First, this method inherently struggles to explore the complex, high-dimensional and multimodal state-action spaces due to its reliance on unimodal Gaussian distribution [Wang2023a]. Second, the frequent penalties imposed by covert constraints further exacerbate this issue by discouraging the agent from effectively leveraging the rare positive reinforcement signals.
To overcome these limitations, we utilize GDM as the policy backbone. Different from fully-connected deep neural networks adopted in the conventional DPG algorithm, GDM generates diverse and high-fidelity action samples through iterative denoising, thereby enabling effective exploration across multimodal regions of the state-action space [Yang2023]. Compared with other generative frameworks, GDM avoids the issue of non-differentiable propagation that arises in generative adversarial networks. Moreover, it also overcomes the limitations of variational autoencoders (VAE), which suffer from unimodal latent priors and susceptibility to posterior collapse [R2-3-1].
@@ 4 $\mathbf{z}\sim\mathscr{N}(\mathbf{0},\mathbf{I})$ if $t>1$ else $\mathbf{z}=\mathbf{0}$;
IV-C2 GDM for Policy Representation
According to [Yang2023], the basic process of GDM includes the forward process and reverse process. Specifically, the forward process of GDM incrementally corrupts a data sample by adding Gaussian noise over a series of time steps, thereby transforming its distribution into one that approximates pure Gaussian noise. More formally, this forward procedure constitutes a Markov chain, which can be defined as follows:
where $T$ is the total number of diffusion steps, and $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathscr{N}(\mathbf{x}_{t};\sqrt{1-\nu_{t}}\mathbf{x}_{t-1},\nu_{t}\mathbf{I})$ represents the transition kernel of the forward process. Moreover, $\nu_{t}\in(0,1)$ is a diffusion weight to control the speed of diffusion. After applying a series of derivations, the joint transition over all time steps can be marginalized to yield a closed-form expression for the distribution of $\mathbf{x}_{t}$ conditioned on $\mathbf{x}_{0}$, which can be computed as follows [Zhang2025]:
where $\mu_{\theta}(\mathbf{x}_{t},t)$ and $\Sigma_{\theta}(\mathbf{x}_{t},t)$ represent the mean and the covariance matrix of the predicting noise at time step $t$, respectively. To incorporate specific semantic attributes into the reverse process, the conditional GDM can provide additional information to guide each denoising step [Dhariwal2021]. Specifically, the reverse transition chain can be further expressed as follows:
Figure 3: The overall architecture of the proposed GDPG algorithm to solve the ASCCOP, where GDM is used to execute the policy representation of the GDPG algorithm and action gradient mechanism is introduced to enable the policy improvement of the GDPG algorithm.
1Initialize GDM network parameters $\theta$, critic network parameters $\phi_{1}$ and $\phi_{2}$ as well as their corresponding target network parameters $t\theta$}, $t\phi_{1}$ and $t\phi_{2}$;
@@ 11         Sample a batch from $\mathcal{D}$;
IV-C3 Action Gradient for Policy Improvement
As shown in Eq. (30), GDM merely learns to approximate the distribution of training samples produced by the current policy $\pi_{\theta}$. As such, it does not intrinsically offer a mechanism to optimize the policy beyond the scope of the data on which it is trained. Different from the conventional DPG algorithm, direct policy improvement through the policy gradient theorem is not feasible in this setting, as the diffusion-based policy is not parameterized straightforwardly but is instead learned via a stochastic process. Therefore, we adopt an action gradient mechanism to enable policy improvement, which is shown in Fig. 2. Specifically, a given state-action pair $(\mathbf{s},\mathbf{a})$ can be converted into a better state-action pair by the gradient ascent method, which can be represented as follows [Zhang2025]:
@@ IV-D Overall Structure and Complexity Analysis
Similar to approaches used in [Fujimoto2018], [Haarnoja2018], we employ a target GDM network $\pi_{t\theta}$, and a double critic networks $Q_{\phi_{1}}$ and $Q_{\phi_{2}}$ alongside their corresponding target critic networks $Q_{t\phi_{1}}$ and $Q_{t\phi_{2}}$ to avoid abrupt distribution shifts and address Q-value overestimation. Specifically, the loss function of the critic network can be expressed as follows:
@@ IV-D1 Training Phase
Let $G$, $B$, $|\bm{\phi}|$, $|\bm{\theta}|$ and $C$ represent the number of training episodes, batch size, the number of critic network parameters, the number of GDM network parameters, and the per-step environment interaction cost, respectively. Then, the total computational complexity during the training phase is $\mathcal{O}(GN(B+C+(4+T)|\bm{\theta}|+6|\bm{\phi}|))$. Specifically, the computational complexity of the training phase is summarized as follows [Zhang2025]:
  • Network Initialization: This process involves initializing two critic networks, a GDM network, and their target networks, thereby increasing the complexity of $\mathcal{O}(4|\bm{\phi}|+2|\bm{\theta}|)$.
  • Action Selection and Environment Interaction: GDM-based action generation at each step needs to performs $T$ denoising steps and interacts with the environment, thereby introducing the complexity of $\mathcal{O}(GN(T|\bm{\theta}|+C))$ during the entire training phase.
  • Network Updates: The overhead of the network update process includes several components. Specifically, the parameter update for the two critic networks incurs the complexity of $\mathcal{O}(2GN|\bm{\phi}|)$. Moreover, action gradient improvements add the complexity of $\mathcal{O}(GNB)$. Additionally, the GDM network update introduces the complexity of $\mathcal{O}(GNT|\bm{\theta}|)$. Finally, the soft update operation results in the complexity of $\mathcal{O}(GN|\bm{\phi}|+2GN|\bm{\theta}|)$. As a result, the overall complexity is $\mathcal{O}(GN(B+2|\bm{\theta}|+4|\bm{\phi}|))$.
@@ IV-D2 Execution Phase
Figure 4: Architecture of GDM network with Mish activation function [Misra2019] for non-monotonic gradient flow, and sinusoidal position embedding for efficient encoding of temporal information in the reverse process.
@@ V-A Simulation Setup
V-A1 Environment, Platform and Model Details
In this work, we consider the service area as a square region with 200 m by 200 m. In this region, 5 ground users are served by an LAP, operating at $H=100$ m, and equipped with a $4\times 4$ active STAR-RIS array to facilitate the wireless covert transmission between a GEO satellite and the ground users. Moreover, the simulations are conducted on a Linux server with an NVIDIA RTX 4090 GPU, using Python 3.9, PyTorch 1.13, and CUDA 11.6. For the proposed GDPG algorithm, we utilize the Adam optimizer with a learning rate $3\times 10^{-4}$ for both the GDM network and the critic networks, while setting $\eta_{a}=3\times 10^{-2}$ for the action gradient. The specific GDM network structure is exhibited in Fig. 4 and other key parameters for the environment and algorithm are summarized in Table II. Note that the parameter setting of the GEO satellite is chosen in accordance with the 3rd Generation Partnership Project (3GPP) Technical Report 38.821 Version 16.2.0 [R1-6-1].
@@ V-A1 Environment, Platform and Model Details
$\sigma_{r}^{2}$, $\sigma_{k}^{2}$, $\sigma_{w}^{2}$ -174 dBm/Hz $\rho$ 3 dB
Algorithm
$\gamma$ 0.95 $T$ 10
@@ V-A2 Benchmarks
  • State-of-the-Art DRL Algorithms: The deep deterministic policy gradient (DDPG) and twin delayed DDPG (TD3) are the earliest actor-critic algorithms designed for continuous action spaces [Fujimoto2018]. Moreover, soft actor-critic (SAC) is based on the maximum entropy reinforcement learning framework and also serves as a baseline [Haarnoja2018]. Besides, we implement a VAE-enabled DPG algorithm by replacing the GDM with the VAE to enable direct assessment of GDM to multimodal exploration and policy stability.
@@
In order to make a fair performance comparison with benchmarks, the neural network architectures of all DRL-based benchmarks comprise two hidden layers for both the actor and critic networks, with each layer containing 256 neurons. Moreover, we follow its original setup by adding target policy smoothing using clipped noise with a standard deviation of 0.2 and a range limit of 0.5 for TD3. For SAC, we enable entropy regularization with an automatic entropy coefficient update. While for DDPG, we add standard Gaussian noise with a standard deviation of 0.1 to each action to enhance exploration. In addition, the learning rates for the actor and critic networks in all DRL-based benchmarks are set to match those of the GDM and critic networks in the proposed GDPG algorithm.
@@ V-B2 Comparison with State-of-the-Art DRL Algorithms
Fig. 6(a) illustrates the reward per episode of the proposed GDPG algorithm in comparison to other state-of-the-art DRL algorithms during the training process. As can be seen, the proposed GDPG algorithm consistently outperforms the other algorithms with a steady increase in reward over the training time. The success of the GDPG algorithm lies in the use of the GDM, which improves the exploration and policy optimization by generating the diverse and high-quality samples through its denoising process, thereby resulting in higher overall rewards, especially in the complex satellite-terrestrial covert communication scenarios. Moreover, DDPG and TD3 demonstrate initial improvements in rewards during the early stages of training, but their progress slows and eventually plateaus. This indicates that although they achieve early gains, their performance is constrained by limited exploration, thereby leading to suboptimal policy. Finally, SAC performs better than DDPG and TD3 with a more consistent increase in rewards. However, it still has lower performance than that of the GDPG algorithm, thereby illustrating that SAC lacks the same level of exploration efficiency and optimization capabilities provided by the proposed GDPG algorithm. In addition, the VAE-enabled DPG algorithm outperforms conventional baselines but underperforms the proposed GDPG algorithm, which confirms that variational autoencoders lack the multimodal expressiveness and stability of GDM in high-dimensional constrained spaces while improving exploration.
Fig. 6(b) shows the mean sum channel capacity per step of the proposed GDPG algorithm in comparison to other state-of-the-art DRL algorithms. The performance achieved by the proposed GDPG algorithm is notably higher than that of other algorithms. This indicates that the GDM-enabled DRL algorithm is more effective in leveraging its representation and exploration capabilities in the large state and action spaces. Another important metric is the fairness index, which is presented in Fig. 6(c). As can be seen, the GDPG and VAE-enabled DPG algorithms maintain the highest fairness indexes across all algorithms. However, the GDPG algorithm achieves the fastest convergence and consistently maintains a higher mean sum channel capacity per step, owing to its denoising process and more effective exploration of the action space.